OpenStudy (anonymous):

if 1/x+1/y=3 and y(4)=4/11, find y'(4) by implicit differentiation. for the answer i got it was 0, i need help

OpenStudy (amistre64):

imply it then :)

OpenStudy (amistre64):

\[Dx(1/x)+Dx(1/y)=Dx(3` )\]

OpenStudy (amistre64):

\[-x'.1/x^2-y'.1/y^2=x'.0\]

OpenStudy (anonymous):

d/dx(1/x)=-1/x^2 and d/dx(1/y)=y'/y^2 1/x^2-y'/y^2

OpenStudy (amistre64):

very good so far :)

OpenStudy (anonymous):

yeah that's what i got then i apply the (4, 4/11) i get -1/16 and -y'/(16/121)

OpenStudy (anonymous):

-1/16-y'/(16/121)

OpenStudy (amistre64):

\[y'=-\frac{y^2}{x^2}=-\left(\frac{y}{x}\right)^2\]

OpenStudy (anonymous):

i got .47265625 in fractions its 121/256 is that right

OpenStudy (amistre64):

\[y' = \frac{16/121}{16/1}=\frac{1}{121}\]right?

OpenStudy (anonymous):

hold on let me do it again

OpenStudy (amistre64):

and i dropped the negative on accident lol

OpenStudy (anonymous):

ohh sorry i flipped the 121and 16 yeah 1/121 is right

OpenStudy (amistre64):

-1/121

OpenStudy (anonymous):

yeah -1/121 is right thanks for the help our professor said it equals 0 on another similar problem in class but the y' formula helps